Suppose a 500. mL flask is filled with 0.70 mol of H2 and 0.60 mol of HI. The following reaction becomes possible: -2H1(g) The equilibrium constant K for this reaction is 2.38 at the temperature of the flask. Calculate the equilibrium molarity of H2. Round your answer to two decimal places.
Answer #1

Ans. Initial [H2] = Moles of H2 / Volume of reaction vessel in L

                        = 0.70 mol / 0.500 L

                        = 1.40 M

Initial [HI] = 0.60 mol / 0.500 L = 1.20 M

# Since there is no I2 gas on the reactant side, the reaction can’t go in forward direction because there is no I2as reactant at all. So, the reaction must go to the backward direction to establish equilibrium.

So, we have to proceed with the reversed reaction        2 HI(g) -----> H2(g) + I2(g)

Reversing a reaction inverses the original value of equilibrium constant.

So, equilibrium constant for 2 HI(g) -----> H2(g) + I2(g)         , K’ =1 / 2.38 = 0.420

# Create an ICE table for the reversed reaction as shown below-

2 HI (g)H2(g) + 12(g) Initial (M) Change (M) Equilibrium (M) 1.2 (-2x) 1.2- 2x 1.4 0 +X

Now,

            K’ = [X (1.4 + X) ] / (1.2 – 2X)2

            Or, 0.420 = (1.4X + X2) / (1.44 + 4X2 – 4.8X)

            Or, 0.420 (1.44 + 4X2 – 4.8X) = 1.4X + X2

            Or, 0.6048 + 1.68X2 – 2.016X = 1.4X + X2

            Or, 0.6048 + 1.68X2 – 2.016X - 1.4X - X2 = 0

            Or, 0.68X2 – 3.416X + 0.6048 = 0

Solving the quadratic equation, we get following two roots-

            X1 = 4.839                 ; X2 = 0.184

Since X can’t be greater than 1.2, reject X1.

Hence, X = 0.184

# Equilibrium [H2] = 1.4 + X = 1.4 + 0.184 = 1.584

            Hence, equilibrium [H2] = 1.58 M